# YUI: Day-ahead Electricity Price Forecasting Using Invariance Simplified Supply and Demand Curve

### v1

@@
Abstract.
In day-ahead electricity market, it is crucial for all market participants to have access to reliable and accurate price forecasts for their decision-making processes. Forecasting methods currently utilized in industrial applications frequently neglect the underlying mechanisms of price formation, while economic research from the perspective of supply and demand have stringent data collection requirements, making it difficult to apply in actual markets. Observing the characteristics of the day-ahead electricity market, we introduce two invariance assumptions to simplify the modeling of supply and demand curves. Upon incorporating the time invariance assumption, we can forecast the supply curve using the market equilibrium points from multiple time slots in the recent period. By introducing the price insensitivity assumption, we can approximate the demand curve using a straight line. The point where these two curves intersect provides us with the forecast price. The proposed model, forecasting supplY and demand cUrve simplified by Invariance, termed as YUI, is more efficient than state-of-the-art methods. Our experiment results in Shanxi day-ahead electricity market show that compared with existing methods, YUI can reduce forecast error by 13.8% in MAE and 28.7% in sMAPE. Code is publicly available at https://github.com/wangln19/YUI.
Electricity price forecasting; Supply and demand curves; Machine learning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 978-1-4503-XXXX-X/18/06isbn: ;ccs: Information systems Data miningccs: Computing methodologies Machine learningccs: Applied computing Forecasting
1. Introduction
Electricity price forecasting is crucial for all market participants. Accurate forecasts are needed for optimal bidding strategies, asset allocation, contract negotiation, risk hedging and facility planning (Li et al., 2007). Electricity plants and consumers can make bidding strategies to avoid economic losses, and power grid corporations can ensure the stability of the grid system operation (Xie et al., 2022a). The dynamics of electricity prices exhibit unique market behaviors, such as unexpected price peaks and price seasonality. This makes accurate price forecasting difficult (Lago et al., 2018b).
Day-ahead electricity market prices, determined by market rules, fluctuate based on the balance between electricity supply and demand. Forecasting electricity prices is a multifaceted field encompassing a broad range of tasks, largely contingent on the specific market in question: Day-Ahead market, Intra-Day markets, or Balancing markets (Maciejowska et al., 2022). Of these, the Day-Ahead market has attracted the most substantial interest (Castelli et al., 2020). In today’s energy sector, the day-ahead electricity market is predominantly characterized by market-based pricing, which is determined by supply and demand, rather than by uniform pricing set by the state or government (Martin et al., 2013).
Currently, methods for day-ahead electricity prices forecasting can be categorized from two distinct perspectives: industrial application and economic research. From the industrial application’s viewpoint, the approach primarily involves the straightforward adaptation of general time-series forecast models to the day-ahead electricity market. Conversely, the economic research’s perspective incorporates market supply and demand dynamics. This approach estimates price by forecasting the intersection point of the supply and demand curves, providing an intuitive understanding of market equilibrium. It also helps to identify factors causing price changes and forecast how individuals and companies will react to these changes.
Existing forecasting methods used in industrial applications often overlook the mechanism of price formation. These methods heavily rely on the incorporation of general machine learning model structures, using historical data from various relevant variables to directly forecast day-ahead prices. Due to their disregard for the economic principles, these methods may overlook certain influential factors, resulting in an accuracy decline (Patel and Shah, 2021).
While economists strive to forecast day-ahead electricity prices using the principle of supply and demand, they encounter challenges in practical applications due to data unavailability (Soloviova and Vargiolu, 2020; Wan et al., 2022). To project the supply curve, it’s essential to gather data on the propensity to produce at distinct price points. Similarly, to project the demand curve, data on the propensity to purchase at specific price points is required. Economists typically assume access to these data, they quantify supply and demand quantity within every price range in the market to construct accurate historical supply and demand curves, subsequently attempting to forecast their temporal evolution. However, procuring such exhaustive data in actual markets is a formidable task because of the private nature of individual buying and selling decisions, which complicates the precise measurement of willingness to buy or sell at each price.
In order to introduce the price mechanism of supply and demand to assist in day-ahead electricity price forecasting under existing data conditions, we propose a viable model forecasting supplY and demand cUrve simplified by Invariance, or YUI for short, to forecast day-ahead electricity price. By examining the traits of the day-ahead electricity market, we introduce two invariance assumptions to streamline the modeling of supply and demand relationship. With the time invariance assumption, we can forecast the supply curve using market equilibrium points from multiple recent time slots, only requiring historical day-ahead prices, supply volumes and capacities data, which are publicly available. By adopting the price insensitivity assumption, we can forecast the demand curve as a straight line parallel to the price-axis, only requiring forecast demand quantities and capacity data, which are also publicly accessible. Through this approach, YUI no longer requires the propensity to produce or purchase data which is hard to get in a changing market. Instead, it leverages data which most markets publicly disclose to forecast supply and demand curves. This method effectively bridges the gap between theoretical economic models and industrial applications. Our YUI framework consists of two main components: forecasting the supply curve and the demand curve. The supply curve is forecast by fitting recent days’ supply curves using historical price equilibrium data, the demand curve is forecast by forecasting the x-intercept of a simplified vertical demand curve. The intersection of the forecast supply and demand curves provides the desired price forecast.
@@ 1. Introduction
  • To the best of our knowledge, this is one of the pioneering efforts toward electricity price forecasting considering the mechanism of price formation practically applied in day-ahead market.
  • We introduce two invariance assumptions from the characteristics of the day-ahead electricity market: time invariance and price insensitivity. We simplify the supply and demand curves forecasting by leveraging these invariance assumptions, especially in the absence of propensity to produce or purchase data.
  • We conduct extensive experiments on two public electricity datasets. The experimental results have demonstrated that our method significantly outperforms existing state-of-the-art baseline methods by $13.8\%$ in MAE and $28.7\%$ in sMAPE.
2. RELATED WORK
Time Series Forecasting. Time series forecasting, a key area in computer science, involves forecasting future values based on past data. It’s characterized by autocorrelation, seasonality, and stationarity. Preprocessing of data is crucial, and various models (Liu et al., 2023a, b; Wu et al., 2022, 2021; Zeng et al., 2023) are used for forecasting, including linear, nonlinear, and deep neural networks. This technique finds wide applications in fields like finance and energy consumption.
Day-ahead Electricity Price Forecasting. Day-ahead electricity price forecasting is vital in the power market, with market participants relying on these forecasts for bidding strategies and risk management. The electricity price exhibits complex characteristics like high volatility and seasonality. Various techniques, including linear regression, ensemble models, and machine learning methods, are employed for forecasting (Uniejewski et al., 2016; Che and Wang, 2010; Wang et al., 2017; Prahara and Hariadi, 2022; Yamin et al., 2004; Lago et al., 2018a; Darudi et al., 2015; Ludwig et al., 2015; Xiong and Qing, 2023; Lago et al., 2021). Each model has its strengths and weaknesses, and combining several models can yield more accurate forecastings. This forecasting plays a crucial role in maximizing economic benefits and mitigating market risks.
Economics Approaches for Electricity Price Forecasting. In the field of economics, the principle of supply and demand is a key technique in price forecasting. This method models the supply and demand curves of auctions. The intersection of these curves, affected by factors like weather conditions and general economy behaviors, determines the market clearing price. There are many research (Wan et al., 2022; Ziel and Steinert, 2016; Pinhão et al., 2022; Soloviova and Vargiolu, 2020) trying to reduce the amount of data and parameters needed, but these methods necessitate historical data from every transaction in the market to model the supply and demand curves, which is challenging to acquire in the day-ahead electricity market.
3. Preliminary and PROBLEM FORMULATION
3.1. The Demand and Supply Model of Microeconomics
In the field of economics, the interplay between supply and demand curves determines the price of a product or service (OpenStax, 2024d). The supply curve represents the quantity that producers are willing to manufacture and sell at various prices. On the other hand, the demand curve represents the quantity that consumers are willing and able to purchase at different prices. The intersection of these two curves, known as the equilibrium, signifies the price at which the quantity supplied equals the quantity demanded (OpenStax, 2024c). Market dynamics typically drive the price towards this equilibrium. In real-world markets, factors such as production costs, consumer preferences, market structures and government regulations influence the equilibrium price by shaping the supply and demand curves (OpenStax, 2024b). In a perfectly competitive market, the cost of the supplier with the highest marginal cost equals the equilibrium price of the market (OpenStax, 2024a). Suppliers must take their costs into account when setting prices. If their price is lower than their marginal cost, they will incur a loss. They must set a price that can at least cover their marginal cost. If they price above marginal cost, consumers will turn to other producers who offer homogeneous products at lower price.
3.2. Day-ahead Electricity Market Structure
Although the regulatory framework for this market differs from country to country, its structure adheres to a standard model: on day $D$, before a specified hour $H$, all market participants must submit their bids for buying or selling energy for each time slot of the target trade day, $D+x(x=1or2)$ (Union, 2024; Haidar et al., 2021). The price for each hourly period is independently determined through an auction-based system (Shi et al., 2023).
Take the market where we deployed, the day-ahead electricity market in Shanxi Province, China, as an example. Here, an authoritative third party considers the anticipated production and costs reported by power producers, forecasts the overall electricity demand quantity, and takes into account various constraints. These constraints encompass grid dispatch limitations, operational characteristics of power plants, and emergency operational reserves. Through complex optimization calculations, the generation schedules of power plants and the dispatch plans of both regional and external power grids are coordinated to ensure that societal electricity demand is met by suppliers. Power generation companies formulate their own production plans based on the finalized schedule from the authoritative third party. The third party adjusts the operational power generation schedule in real time during the trading day to account for any fluctuations in electricity demand and supply. After the trading day, the third party compensates the cost of power generation according to the price settlement rule specified in the operation rules (sha, 2023). Through this pricing process of balancing supply and demand, the day-ahead electricity price at the $h$ moment on $D+x$ day in the entire Shanxi market is determined.
3.3. Problem Formulation
@@
Our objective is to forecast every time slot’s day-ahead electricity price on day $D+x$, prior to the market closure on day $D$. The variables at our disposal encompass data on day-ahead electricity prices, available generation capacity and market demand quantity. The available generation capacity represents the maximum power that a generator set can produce under its current operating conditions, serving as an indicator of the power generators’ production capacity and scale. The load rate is obtained by dividing the quantity of supply by the available generation capacity. It provides insight into the operational status of power generation equipment. Suppose that $P_{h}^{D}$ is the day-ahead electricity price at the $h$ moment on $D$ day, $C^{D}$ is the available generation capacity on $D$ day, $Qs_{h}^{D}$ is the supply quantity at the $h$ moment on $D$ day, $Qd_{h}^{D}$ is the market demand quantity at the $h$ moment on $D$ day and $LR_{h}^{D}$ is the load rate at the $h$ moment on $D$ day.
To ensure transparent management and adequate supply, the authoritative third parties typically disclose operational data on day $D$, including the historical data of these variables and the forecast values of some variables on day $D+x$. Suppose n is the number of trading time slot on day $D+x$, on trading day $D$, we can obtain forecast values for weather conditions such as temperature and wind speed, note as $\hat{Wd}_{h}^{D+x}(h=1,...,n)$, as well as the forecast market demand quantities published by third parties, note as $\hat{Qd}_{h}^{D+x}(h=1,...,n)$.
4. Invariance assumptions for modeling the supply-demand curve
4.1. Challenges of Modeling Supply and Demand Curves
The proposed model, which uses supply and demand curves to forecast day-ahead electricity prices, necessitates complete production and purchasing willingness data over a period of time. This data should encompass the production willingness of suppliers as well as the purchasing willingness of buyers across all price segments. These models are designed to quantify supply and demand within a specific price range in the market (Pinhão et al., 2022; Soloviova and Vargiolu, 2020; Wan et al., 2022; Ziel and Steinert, 2016). After aggregating all the data, they derive the actual supply and demand curves and proceed to fit them. This process necessitates rigorous data collection, as any noise or missing data can significantly affect the accuracy of the curve fitting.
The inability of many day-ahead markets to gather comprehensive willingness information renders these methods impractical. Buyers and sellers often choose to keep their willingness to buy or sell at various price points confidential (Zhao et al., 2021; Hosseini Imani et al., 2023). This reluctance stems from the sensitivity of such information, which could potentially reveal their strategies and negotiation power. Moreover, the willingness to buy or sell is not a static factor (Zhang and Chang, 2021; Dowling, 1995). It can fluctuate rapidly due to a variety of influences such as market trends, economic indicators, political events, and technological advancements.
4.2. Invariance Assumptions in Day-Ahead Electricity Market
To tackle the challenge of data insufficiency in forecasting supply and demand curves, we introduce two invariance assumptions.
Figure 1. Invariance Diagram: The supply curve, represented in blue, consistently maintains its position across different time points, illustrating time invariance. The demand curve, depicted in green, remains stable within a specific price range, demonstrating price insensitivity.
[…]
Invariance Diagram
Assumption 1: Temporal Invariance of the Supply Load Curve. In a highly competitive environment such as the day-ahead electricity market, it is reasonable to consider the supply curve based on the load rate, as remaining constant over a specific time period. Commodities offered by different power generators are entirely homogenized, resulting in the supply curve equaling to the marginal cost (Roberts, 1987). It’s important to note that within a supplier, the marginal cost is primarily dictated by the load rate instead of the supply quantity. This phenomenon is closely tied to the operational characteristics of the generator (Roozbehani et al., 2022): costs increase rapidly as power output rises when generating unit transitioning from standby to startup, slow down upon reaching a stable segment. If supply exceeds the existing unit’s upper limit, the activation of a backup unit is necessitated, leading to another costs surge. Owing to managerial inertia, power plants typically do not alter the operational units over several consecutive days (Dow et al., 2018). This leads to a relatively stable relationship between the marginal cost and the load rate during this period. Ultimately, this is manifested in the following way: using the load rate instead of supply quantity as the horizontal axis of the supply curve, the relationship between the price (represented on the vertical axis) and the load factor exhibits time invariance over recent days. The converting is rather easy, by dividing the supply quantity by the available generation capacity on that day. As Fig. 1 shows, the supply curve, represented in blue, consistently maintains its position across different time points, illustrating time invariance.
As depicted in Fig. 2, we choose a time period from March 23, 2023, to April 3, 2023, within the Shanxi day-ahead electricity market to illustrate the temporal invariance. Fig. 2(a) reveals that the shape of the supply curve remains relatively stable across several consecutive days, necessitating only minor shifts to the left or right for an approximate overlap. By normalizing the absolute value of the supply by the total available generation capacity for the day, or in other words, using the load rate instead of the supply quantity, we can align the supply curves of adjacent dates. This is demonstrated in Fig. 2(b). This alignment uncovers the temporal invariance of the supply curve over a series of consecutive days. Beside, as shown in Fig. 2(c), when there is a long time interval between two dates, the similarity of their supply curves will decrease.
(a) The supply curves, where the x-axis represents supply quantity, retains its form over successive days, necessitating only minor shifts to the left or right for congruence.
@@ 4.2. Invariance Assumptions in Day-Ahead Electricity Market
(b) The aligned supply curves show temporal invariance after normalizing the supply by capacity, where the x-axis represents the load rate.
@@
(c) The disparity in supply curves intensifies when comparing dates that are spaced apart by substantial intervals.
@@
Figure 2. The supply curves of the Shanxi market show temporal invariance within a period of adjacent dates.
Assumption 2: Price Insensitivity of the Demand Curve. Current research on the electricity market is in its nascent stages, with a lack of theoretical studies on the characteristics of the demand curve. We have observed that the quantity of electricity demand is highly inelastic within a reasonable price range and is minimally affected by price fluctuations. To simplify, we have proposed the assumption of price insensitivity. This phenomenon is widely observed, the same-year price elasticity of state-level electricity demand in the U.S. is small, around 0.10.1–0.1– 0.1 for all sectors (Burke and Abayasekara, 2018). There are many explanations for this phenomenon. Electricity, being an essential energy source for modern life and industrial production, has a highly inelastic demand quantity (Zohuri, 2016; Vaghela et al., 2023). Increases in electricity prices have minimal impact on consumption levels due to the difficulty consumers and businesses face in reducing their electricity usage (Papouis et al., 2023). Besides, current electricity storage technology is still underdeveloped, preventing consumers from circumventing price fluctuations through electricity storage (Initiative, 2023). Furthermore, in many regions, electricity prices lack full transparency, and price information is often inaccessible to consumers in time, inhibiting their ability to adjust electricity usage based on price fluctuations (Xu et al., [n. d.]). As a result, the influence of price factors on electricity demand quantity in the electricity market is minimal, manifesting in the demand curve has a large slope within a reasonable price range, can be approximated as a line parallel to the price-axis and unrelated to price. As Fig. 1 shows, the demand curve in green remains stable within a specific price range, demonstrating price insensitivity.
We can verify the price insensitivity of the demand curve from the following perspective: In the electricity market, estimates of future demand are very common. We usually observe that the gap between the forecast demand and the day-ahead market demand is small, and the difference between the forecast demand and the actual demand has almost no relation to the day-ahead price(According to general demand theory, the value of forecast demand minus actual demand should be positively correlated with price). To support this observation, we analyzed two datasets. In the ISO dataset, the Spearman’s correlation coefficient is 0.108 and the Pearson’s correlation coefficient is 0.039, indicating a weak positive correlation. However, in the Shanxi dataset, both the Spearman’s correlation and the Pearson’s correlation are negative, -0.242 and -0.243 respectively, suggesting a weak negative correlation. These findings suggest that the demand curve in the electricity market exhibits price insensitivity.
4.3. Refining Supply and Demand Curve forecasting Through Invariances
With the introduction of the aforementioned invariances, we can simplify the original supply and demand curve.
Figure 3. Numerous historical equilibrium points shape a supply curve
[…]
The supply curve intersects with these demand curves at numerous equilibrium points.
Forecasting the supply curve using the recent supply curve. The temporal invariance of the supply curve allows for forecasting future supply curves using past equilibrium points data to approximate the day’s supply curve. Our challenge now is to ascertain the recent supply curve without access to data of willingness to buy or sell at each price. It is understood that there are many trading time slots within a day. At each of these slots, we can observe the equilibrium day-ahead electricity price and power generation across the entire market, which correspond to the coordinates of the intersection in the supply and demand curve model. While the position of the demand curve varies significantly within a day, the supply curve remains relatively stable. The different transaction prices and volumes at different times are caused by this supply curve intersecting with different demand curves. These intersections are all on the same supply curve, so these intersections can be used to fit the supply curve in reverse. We can observe these equilibrium points and use them to fit a supply curve that approximates the day’s supply curve, as Fig. 3 shows.
Forecasting the demand curve with the forecast load rate. The price insensitivity of the demand curve informs us that the demand curve in the day-ahead electricity market can be approximated as a line parallel to the price-axis and unrelated to price within a reasonable price range. This line parallel to the price-axis necessitates only one parameter - the market’s demand quantity at the target moment. Conveniently, many third parties currently disclose the market demand quantity forecasting to ensure that societal electricity demand quantity can be fully met, especially during peak periods. However, the supply curve we employ uses the load rate as the x-axis. To synchronize the coordinates of both the supply and demand curves, in addition to the forecast market demand quantity, we need to forecast the total available generation capacity of the supply side on the target date, and obtain the load rate by dividing the demand quantity by the capacity. The available generation capacity on the target date is influenced by variables such as weather and forecast demand quantity (Titus et al., 2020). This is a typical time series forecasting problem of using multiple variables to forecast a single variable.
5. METHODOLOGY
5.1. Framework Overview
Figure 4. Overview of YUI framework.
[…]
Overview of YUI framework
Our proposed framework, referred to as YUI, is composed of two primary components: forecasting supply curve and forecasting demand curve(as shown in Fig.  4).
In the supply curve forecasting component, we aim to forecast the supply curve by fitting it to recent days’ data. We select the most suitable historical data from dates with similar curves. Fitting a supply curve using historical price equilibrium data is an optimization problem, as shown in Fig. 3. The goal is to identify a supply curve in the price-load rate plane that minimizes the distance to a known set of dispersed points. Each point represents a trading time slot from recent days, with the x-coordinate being the load rate and the y-coordinate being the day-ahead electricity price for that time slot. These points come from the intersections of the supply curve with various demand curves, and all this data can be publicly obtained from the market.
In the demand curve forecasting component, we forecast the demand curve by forecasting the load rate. We use the forecast demand quantity, usually provided by a third party, and forecast weather data to forecast the available generation capacity for the target date. This forecast is based on the observed relationship between weather, demand quantity, and capacity. At the same time, we generate another forecast for the available generation capacity of the target date, this time based on the time series characteristics of the capacity itself. By combining these two forecasts, we can achieve a relatively accurate forecast. The x-intercept for the approximate demand curve can then be determined by dividing the forecast demand quantity by the forecast available generation capacity.
Finally, we draw the result curves from both components together. The intersection of the forecast supply and demand curves yields the price forecasting value we seek. The data we use are available in the electricity market.
5.2. Forecasting the Supply Curve
In the component for forecasting the supply curve, our goal is to forecast the supply curve by fitting it to the data from recent days.

### v2

# Revisiting Day-ahead Electricity Price: Simple Model Save Millions

@@
Abstract
Accurate day-ahead electricity price forecasting is essential for residential welfare, yet current methods often fall short in forecast accuracy. We observe that commonly used time series models struggle to utilize the prior correlation between price and demand-supply, which, we found, can contribute a lot to a reliable electricity price forecaster. Leveraging this prior, we propose a simple piecewise linear model that significantly enhances forecast accuracy by directly deriving prices from readily forecastable demand-supply values. Experiments in the day-ahead electricity markets of Shanxi province and ISO New England reveal that such forecasts could potentially save residents millions of dollars a year compared to existing methods. Our findings underscore the value of suitably integrating time series modeling with economic prior for enhanced electricity price forecasting accuracy.
Introduction
Electricity price forecasting is closely linked to residential welfare and deserves thorough investigation. Modern society relies heavily on a stable supply of electrical energy, but the cost can be a significant financial burden. Ensuring reliable power supply while reducing living costs is vital, especially for low-income populations. In some countries, residents buy electricity directly from the market, while in others, state-owned companies purchase and retail it. Regardless of the form, the electricity price in the power market directly affects residents’ cost. Prices fluctuate throughout the day, leading to varying consumption costs (Bielen et al. 2017). Advanced energy storage systems allow residents to buy electricity during low-price periods and use stored energy during high-price periods, reducing expenses. This strategy relies on accurate price forecasts, which directly impact financial savings (Zheng et al. 2020; Gil et al. 2012). Therefore, effective electricity price forecasting is essential for optimizing residents’ revenue and enhancing societal welfare.
The day-ahead market is a pivotal part of the electricity market, and this paper focuses on forecasting day-ahead electricity prices. Forecasting electricity prices involves various tasks depending on the specific market: Day-ahead, Intra-day, or Balancing markets (Maciejowska, Uniejewski, and Weron 2022). Among these, the day-ahead market is the most significant, attracting the majority of traders (Castelli, Groznik, and Popovič 2020).
Current methods for forecasting day-ahead electricity prices often underperform. These approaches primarily involve the straightforward adaptation of general Time Series Forecast (TSF) models to the day-ahead electricity market (Lago et al. 2021). However, the electricity market is characterized by large price fluctuations, a lack of clear periodicity and temporal non-stationarity. Consequently, existing methods may not be effective (Patel and Shah 2021).
Current forecasting methods are limited by their inability to capture the strong, variable correlations between prices and supply-demand. The price formation mechanism in the electricity market is clear and well-understood; economics indicates that prices are determined by the balance between supply and demand (Martin, Müller, and Pokutta 2013). Therefore, the correlation between electricity prices and market supply-demand is crucial for accurate forecasting (Ziel and Steinert 2016). However, TSF methods struggle with this. Some approaches just ignore the impact of supply-demand variables (Zeng et al. 2023; Nie et al. 2022), while others trying to use data-driven methods to model these correlations (Liu et al. 2023a; Zhang and Yan 2023). Nonetheless, we find that in trained models, these correlations do not significantly contribute to forecasting results.
We propose a simple model to forecast day-ahead electricity prices by modeling the correlation between prices and supply-demand. Instead of directly using TSF models to forecast prices, we utilize supply-related variables forecasted by TSF models and focus on establishing the correlation between these variables and prices based on the price formation mechanism. This approach avoids the weaknesses of current TSF models in capturing variable correlations and provides better interpretability. Specifically, we introduce the assumption of the short-term stability of the supply curve, under which we can forecast the correlation between supply and price using recent historical data. We fit the supply curve using a simple Correlation-based Piecewise Linear Model (CoPiLinear), then use this and the forecasted demand to forecast the price, as Figure 1 shows.
Figure 1: Overview of CoPiLinear.
@@ Introduction
  • By analyzing the price formation mechanism in the market, we identify the limitations of current models. We introduced a new correlation-based forecasting approach, bridging the gap between the day-ahead electricity price forecasting problem and existing TSF models.
  • To our knowledge, this is one of the first efforts to forecast electricity prices by considering the price formation mechanism in the day-ahead market. We introduce a short-term stability assumption based on market characteristics to simplify supply curve forecasting.
  • Extensive experiments show that our simple method outperforms complex state-of-the-art time (SOTA) models. We demonstrate the revenue of improved electricity price forecasting accuracy for residents, highlighting the significant potential and social impact of the day-ahead electricity price forecasting problem.
Related Work
Recent studies on day-ahead electricity price forecasting have employed models directly migrated from existing TSF methods, can be categorized into two types based on the selection of known variables. The first type focuses on the temporal dependency of the target variable, capturing patterns like periodicity and trends. This includes univariate and channel-independent multivariate TSF models. The second type, in addition to considering the temporal dependency of the target variable, also takes into account the correlation between the target variable and auxiliary variables, represented by various channel-dependent multivariate TSF models.
TSF models emphasize end-to-end forecasting. Advanced second type methods generally capture inter-variable correlations through modules like MLP and attention in parts of the model’s substructure. This approach introduces additional parameters and makes it challenging to intuitively explain the correlations between variables. Moreover, advanced second type methods do not achieve more accurate forecasts with more information as intended. This has been demonstrated by the superior performance of first-type methods like TimesNet (Wu et al. 2022) and PatchTST (Nie et al. 2022) on numerous multivariate datasets.
Modeling the price formation mechanism aligns with economic theory but faces practical challenges due to data unavailability. Accurate projection of supply and demand curves requires data on production and purchase propensities at various price points. However, obtaining comprehensive data is difficult in extensive markets due to the private nature of individual buying and selling decisions, complicating precise measurements of willingness to buy or sell at each price.
Reference Model Domain-Specific Variable Correlation
(Zhao et al. 2017) SARIMA T F
(McHugh et al. 2019) SARIMA T F
(Xiong and Qing 2023) VMD-LSTM T F
(Zeng et al. 2023) DLinear F F
(Wu et al. 2022) TimesNet F F
(Nie et al. 2022) PatchTST F F
(Uniejewski, Nowotarski, and Weron 2016) Linear T T
(Che and Wang 2010) SVM T T
(Wang et al. 2017) SVM T T
(Prahara and Hariadi 2022) SVM T T
(Manfre Jaimes et al. 2023) XGBoost & LSTM T T
(Xie et al. 2022) XGBoost T T
(Yamin, Shahidehpour, and Li 2004) DNN T T
(Darudi, Bashari, and Javidi 2015) DNN T T
(Ludwig, Feuerriegel, and Neumann 2015) Lasso-RF T T
(Liu et al. 2023b) Koopa F T
(Chen et al. 2023b) TSMixer F T
(Zhou et al. 2021) Informer F T
(Wu et al. 2021) Autoformer F T
(Zhou et al. 2022) FEDformer F T
(Zhang and Yan 2023) Crossformer F T
(Liu et al. 2023a) iTransformer F T
(Ziel and Steinert 2016) Supply-Demand T
(Soloviova and Vargiolu 2020) Supply-Demand T
(Wan, Kober, and Densing 2022) Supply-Demand T
(Pinhão, Fonseca, and Covas 2022) Supply-Demand T
Table 1: Summary of related work.
The summary of related work is listed in Table 1. The Model column indicates the primary model used for forecasting in the literature. The Domain-Specific column shows whether the method is tailored for day-ahead electricity price forecasting. A $\bullet$ in the Domain-Specific column denotes methods developed for economic research, not applicable in many markets. The Variable Correlation column reflects if the literature leverages variable correlations for forecasting.
Preliminary and Problem Formulation
Figure 2: Shanxi market transaction process.
Day-ahead Electricity Market Structure
While regulatory frameworks for day-ahead electricity markets vary by country, their structure generally follows a standard model (Union 2024; Haidar, Carmen, and Muttaqi 2021). Prices for each time slot is independently determined through an auction-based system (Shi et al. 2023), driven by supply and demand dynamics. Take the day-ahead electricity market in Shanxi Province, China, as an example. There are 96 trading time slots in a day, each 15 minutes long. As shown in Figure 2, an authoritative third party announces the day-ahead electricity prices for day $D+1$ and other forecast variables on day D, then market participants submit their bids before the day $D+2$ market closes. The third party considers anticipated production and costs reported by power producers and accounts for constraints such as grid dispatch limitations, operational characteristics of power plants, and emergency reserves. Through optimization calculations, generation schedules and dispatch plans are coordinated to meet societal electricity demand. Power generation companies then formulate their production plans based on the finalized schedule (sha 2023). Through this pricing process of balancing supply and demand, the day-ahead electricity prices for day $D+2$ in the entire Shanxi market is determined.
Limitation of Time Series Forecasting Models
On July 17, 2023 (highlighted in yellow), the forecasted demand peaked for the week, coinciding with an exceptionally high day-ahead electricity price.
Forecasting models that rely solely on temporal dependence struggle to accurately forecast high day-ahead electricity prices, as seen on July 17, 2023.
Models that capture variable correlations show relative success in forecasting price increase trends, though the magnitude of these increases is limited.
Figure 3: Actual and forecasted prices of ISO New England on July 17, 2023.
Day-ahead electricity prices show significant dispersion in the high-price range. In the ISO New England market, while 75% of prices are below \$40/MWh, peaks can reach \$300/MWh. This irregularity and lack of clear time series patterns result in poor performance of baseline models, especially time series forecasting models, in forecasting high prices (Rahaman et al. 2019; Xu 2020).
High prices are mainly driven by high demand, which can help forecast these irregularities. However, TSF models struggle to capture the correlation between prices and supply-demand dynamics. According to the price formation mechanism, forecasted demand should significantly impacts forecasted price. Yet, even with substantial changes in demands, existing models show minimal price variation. This inconsistency with real market behavior indicates that these models do not adequately incorporate supply and demand variables, leading to poor forecasting performance. See Appendix A for detailed experiment results.
In Figure 3, the ISO New England Dataset from July 17, 2023, shows forecasted peak demand surpassing the previous week’s peak, indicating high prices. Models based solely on temporal dependence fail to account for forecasted demand, leading to inaccurate high price forecasts. While existing TSF models that capture variable correlations perform better in predicting price trends, they underestimate price increases during significant demand surges, overlooking the price formation mechanism.
Problem Formulation
@@
Our goal is to forecast the day-ahead electricity price for each time slot on day $D+x$ before market closure. The variables at our disposal include day-ahead electricity prices, capacity, weather data and market demand. The capacity indicates the maximum power the generator can produce. The load rate, calculated by dividing supply quantity by capacity, shows the operational status of power generation equipment.
Suppose that $P_{h}^{D}$ is the day-ahead electricity price at the $h$ moment on $D$ day, $C^{D}$ is the capacity on $D$ day, $Qs_{h}^{D}$ is the supply quantity at the $h$ moment on $D$ day, $Qd_{h}^{D}$ is the market demand quantity at the $h$ moment on $D$ day and $LR_{h}^{D}$ is the load rate at the $h$ moment on $D$ day. The forecast weather data for day $D+x$ include temperature and wind speed, denoted as $\hat{W}_{h}^{D+x}$, and the forecast market demand quantities published by third parties, denoted as $\hat{Qd}_{h}^{D+x}$.
This study evaluates the economic revenue of day-ahead price forecasts for residents. Based on related work (Zheng et al. 2020), we formulate an optimization problem for the purchasing strategy, using given forecast prices and the energy storage systems. The cost savings from this strategy represent the economic revenue of price forecasts, and improved forecast accuracy enhances the revenue. The purchasing cost $C$ reaches its theoretical minimum $C_{a}$ if forecasts match actual prices perfectly. Without the strategy, the cost is $C_{0}$. The economic revenue $R$ equals $C_{0}-C$. To assess the impact of forecast accuracy, we calculate $\alpha=\frac{R}{C_{0}-C_{a}}$, reflecting the percentage of the maximum theoretical revenue achieved by the current forecast.
Methodology
Framework Overview
We notice that the day-ahead electricity price is determined by the supply and demand, so we directly use this correlation to forecast the price. Forecasting precise supply and demand curves is impractical due to the need to forecast individual trading decisions at various price levels. Obtaining such detailed data is challenging because of the private nature of these decisions.
To address this issue, we introduce the assumption of the short-term stability of the supply curve. Using recent historical data, we reconstruct the supply-price correlation and obtain a forecast for the future supply curve. We fit the supply-price relationship using a simple Correlation-based Piecewise Linear Model (CoPiLinear). With this curve and supply volume forecasted by TSF models, we can forecast the price, as Figure 1 shows.
Short-term Stability Assumptions
The supply curves, plotted with supply quantity on the x-axis, retain their shape over successive days, needing only slight left or right adjustments for alignment.
@@
After normalizing supply by capacity, the aligned supply curves demonstrate short-term stability, with the x-axis representing the load rate.
@@
The disparity in supply curves increases when comparing dates separated by longer intervals.
@@
Figure 4: The supply curves of the Shanxi market show short-term stability within a period of adjacent dates.
It is reasonable to assume the supply curve remains constant over a specific period. Commodities from different electricity power generators are homogenized, making the supply curve equivalent to the marginal cost (Roberts 1987). Notably, within a supplier, the marginal cost is primarily dictated by the load rate rather than the supply quantity. The load rate reflects the generator’s operating conditions, which directly impact costs (Roozbehani et al. 2022): costs rise rapidly when transitioning from standby to startup, slow down upon reaching a stable segment, and surge again if a backup unit is activated. Due to managerial inertia, power plants typically do not alter operational units over several consecutive days, leading to a stable correlation between marginal cost and load rate during this period. More details about the short-term stability are listed in the Appendix B.
Using the load rate instead of supply quantity as the horizontal axis of the supply curve, the correlation between price and load factor exhibits short-term stability over recent days. This conversion is straightforward by dividing the supply quantity by the capacity on that day.
As shown in Figure 4, we selected a period from March 23, 2023, to April 3, 2023, within the Shanxi day-ahead electricity market to illustrate the short-term stability. The supply curve shape remains relatively stable across several consecutive days, requiring only minor shifts for approximate overlap. By using the load rate instead of supply quantity, we can align the supply curves of adjacent dates. This alignment reveals the short-term stability of the supply curve over consecutive days. Additionally, the similarity of supply curves decreases with longer intervals between dates.
Fitting Supply Curve by CoPiLinear
Figure 5: Left Image: The green curve shows varying demand over time, while the blue curve represents a stable supply. Their intersection points indicate market prices. Right Image: The scatter points represent supply-demand intersections over time, from which a supply curve can be derived.
Assuming short-term stability, we can forecast the supply curve by fitting it to recent days’ data. Our challenge is to determine the recent supply curve without data on willingness to buy or sell at each price. While the demand curve position varies significantly at different time slots, the supply curve remains relatively stable. We can get historical equilibrium day-ahead electricity prices and power supply volume at various trading time slots, corresponding to the intersections of the supply and demand curve. These intersections, all on the same supply curve, can be used to fit the supply curve in reverse, as shown in Figure 5.